# AG-SLAM: Active Gaussian Splatting SLAM

### v1

@@ Appendix A Appendix
A.1 Implementation Details
The 2D occupancy map’s resolution is 5cm. For each single frontier pixel on the 2D map, we add 200 3D Gaussians, which are uniformly distributed in the 3D cube above it. Other parameters like color, opacity, and scale are generated uniformly between 0 and 1. When there are frontiers on 2D map, we choose the next frontier to be explored by the area of each frontier divided by the distance. When no frontier exists, we select the top 20% of Gaussians with the highest score. These Gaussians are grouped using DBSCAN [9]. The largest cluster is selected for candidate pose generation. Candidates are uniformly sampled in the range between 0.3m to 1m, facing towards the selected position. Only the poses in free space are kept for path-level selection. The importance factor $\eta$ in Eq. 17 is set to 5 across all experiments. The source code for this project will be made public no later than the publication of this paper.
@@ A.1 Implementation Details
A.2 Results for Each Scene in Gibson and HM3D Dataset
@@
AG-SLAM Refer to caption Refer to caption Refer to caption

### v2

@@ Appendix A Appendix
In this supplementary material, we discuss things we left over in our main paper due to page constraints. We provide more background information about 3DGS and the derivative with respect to localization parameters in Sec. A.1. We discussed more implementation details in Sec. A.2 and provided the prompt we use for the Multimodal LLM and an example of the interaction in Sec. A.5 to help reproduce our results. The source code of this project will be made public soon. We also include detailed quantitative and qualitative results in Sec. A.3 and Sec. A.4.
A.1 Additional Background Information
For the completeness of our method, we also provide the key definition for the 3D Gaussian Splatting backbone [33] and 3D Gaussian SLAM [44]. In 3D Gaussian Splatting [33], the rendered pixel color is calculated by composing all 3D Gaussians projected in a tile.
$\hat{C}(\text{r})=\sum_{i=1}^{N_{s}}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right){\bf c}_{i}$ (14)
$\alpha_{i}=\text{exp}\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)(1-\text{exp}(-\sigma_{i}\delta_{i}))$ (15)
$\delta_{i}=t_{i+1}-t_{i}$ represents the distance between adjacent samples, and $N_{s}$ indicates the number of samples. ${\bf c}_{i}$ is the color of each 3D Gaussian given the current view direction ${\bf d}$ and $\sigma_{i}$ is given by evaluating a 2D Gaussian with covariance $\Sigma$.
The Jacobian of the localization parameters are defined as:
$\frac{\mathcal{D}{\boldsymbol{\mu}_{C}}}{\mathcal{D}{{\bf x}}}=\begin{bmatrix}\boldsymbol{I}&-\boldsymbol{\mu}_{C}^{\times}\end{bmatrix}\text{ and }\frac{\mathcal{D}{\mathbf{W}}}{\mathcal{D}{{\bf x}}}=\begin{bmatrix}\mathbf{0}&-\mathbf{W}_{:,1}^{\times}\\ \mathbf{0}&-\mathbf{W}_{:,2}^{\times}\\ \mathbf{0}&-\mathbf{W}_{:,3}^{\times}\\ \end{bmatrix}~{},$ (16)
where × denotes the skew symmetric matrix of a 3D vector, and $\mathbf{W}_{:,i}$ refers to the $i$th column of the matrix.
Unlike PSNR, the EIG can be computed without ground truth images, making it possible to perform view selection during exploration.
A.2 Implementation Details
The 2D occupancy map’s resolution is 5cm. For each single frontier pixel on the 2D map, we add 200 3D Gaussians, which are uniformly distributed in the 3D cube above it. Other parameters like color, opacity, and scale are generated uniformly between 0 and 1. When there are frontiers on the 2D map, we choose the next frontier to be explored by the area of each frontier divided by the distance. When no frontier exists, we select the top 20% of Gaussians with the highest score. These Gaussians are grouped using DBSCAN [19]. The largest cluster is selected for candidate pose generation. Candidates are uniformly sampled in the range between 0.3m to 1m, facing towards the selected position. Only the poses in free space are kept for path-level selection. The importance factor $\eta$ in Eq. 13 is set to 5 across all experiments. The source code for this project will be made public no later than the publication of this paper.
@@ A.2 Implementation Details
The compute of Fisher Information in our experiments takes 30 frames per second on our Nvidia L40 GPU. The time on path planning will not significantly impact the robot’s motion if the path is planned ahead of time and runs in parallel as the robot moves. The regularizer $\lambda$ in Eq. 4 is $10^{-6}$ across all experiments. The weight parameters $\eta$ in Eq. 13 is 0.2 across all experiments.
A.3 Results for Each Scene in Gibson and HM3D Dataset
@@
A.4 Detailed version of qualitative results
We provide larger versions of the qualitative rendering comparisons from the main paper, Fig. 11 shows the Gibson scenes and Fig. 12 shows the HM3D scenes.
A.5 Example of Using Multimodal LLM
We provide an example of our interaction with the multimodal LLM in Fig 13.
1
2<system> You are an AI assistant that can analyze images and plan a long-term goal for the exploration task of a ground robot.
3You will be given a bird-eye view image of a scene.
4The goal is to plan a long-term exploration mission for a robot to traverse the area.
5The robot's task is to explore the terrain efficiently, identifying important areas, potential obstacles, and unvisited areas.
6Please analyze the image and select a long-term goal from the candidates for the robot to explore the area.
7Empty space doesn't always mean they are unvisited regions, sometimes it's just outside the floor plan of this scene.
8We are allowed to explore a total of <TOTAL\_STEPS> steps and this is step <STEP\_ID>.
9Therefore, it's better to select a space that is close to the visited regions but still unvisited and not behind the walls.
10The current location of the robot is marked with the blue star(*) marker.
11The last frontier you selected is marked with a yellow diamond shape.
12The visited path is painted as green lines in the image.
13Note that you don't have to select the closest point to the robot, but the point that is most likely to be unvisited and important to explore.
14As you can see, there are <NUM\_FRAME> candidate points to select from.
15They are numbered from 0 to <NUM\_FRAME - 1> in red color.
16If you find all the goals are not necessary to explore and we should instead focus on improving existing reconstruction, please give -1 in the `target` entry of the JSON.
17Please provide a detailed exploration plan and select an exploration target with reasons in the JSON format as shown below.
18
19
20```
21{
22"target": 2, "reason": "The target is located at an unvisited region of the image and seems to be an unvisited bedroom"
23}
24```
25Do not cut off the JSON and generate the full JSON.
26</system>
27
28
29<user>: I have a bird-eye view image of a scene. The goal is to plan a long-term exploration mission for a robot to traverse the area. Please analyze the attached image and provide the exploration plan first and then an exploration target in the specified JSON format.
30</user>
@@ A.5 Example of Using Multimodal LLM
Ours Refer to caption Refer to caption Refer to caption
@@
UPEN Refer to caption Refer to caption Refer to caption
FBE Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure 10: Test Rendering Qualitative Visualization on Habitat-Matterport 3D Dataset All the renderings are from the test view of the Habitat-Matterport 3D dataset.
ANS [7] Active-INR [85] UPEN [22]
Refer to caption Refer to caption Refer to caption
ExplORB [53] FBE [82] Ours
Refer to caption Refer to caption Refer to caption
ANS [7] Active-INR [85] UPEN [22]
Refer to caption Refer to caption Refer to caption
ExplORB [53] FBE [82] Ours
Refer to caption Refer to caption Refer to caption
UPEN [22] FBE [82] Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: An Example about a Dialog with the Multi-modal LLM. The chat box on the top is the question from our method, and the bottom is the response from GPT-4o, which we parsed to get a long-term goal for our path proposal. We provided the contextual information in both the textual prompt and the image.